# Adversarial Coevolutionary Illumination with Generational Adversarial MAP-Elites

### v1

@@ Introduction
  • GAME-MO (no VEM): full variant with a 2D handcrafted behavior space equal to the evaluated side average remaining health and the time before completion (i.e., one side wins or $100$ steps timeout) instead of a VEM;
  • Quality-only: MT-ME instead of MT-MB ME, i.e., each task has only one solution;
  • Diversity-only: do not use fitness, only rule (e.1).
We run three replications of $20$ generations for each variant with 100 000100000100\,000100 000 evaluations of $100$ steps per generation, $100$ tasks per generation, and $25$ cells per task’s archive.
@@
After each run, we realize a tournament between each generation’s tasks, $10$ Blue generations against $10$ Red generations (i.e., 1 00010001\,0001 000 Red elites against 1 00010001\,0001 000 Blue elites), resulting in 1 000 00010000001\,000\,0001 000 000 evaluations. This allows us to study whether there are intergenerational improvements.
@@
To measure the variations of solution size, we measure the number of nodes of the BT elites (Fig. 3.a), and for behavioral complexity, we measure the entropy of the distribution of actions performed over time and units of the same side (Fig. 3.b).
Visual Diversity and Quality Visual Diversity
Coverage and QD-Score are classic measures of QD algorithms, but are not easily computed using the archive in a large behavior space. As a proxy, we compute a single 2D projection using PCA (the two components capturing $22$% and $12$% of the explained variance) for all the behaviors of the intergenerational tournaments for all replications (Fig. 4), discretize this space with a $100\times 100$ grid, and compute the corresponding elites. The Coverage (Fig. 3.c) corresponds to the proportion of non-empty cells, and the QD-Score (Fig. 3.d) to the average fitness of the elites.
@@
One limitation with computing the quality directly from the fitness in the QD-Score is that the fitness can be high because the opposite side is bad, and not because the winning side is good. To measure a less ambiguous quality of the elites found by each method, we follow Dharna et al., (2024) and select the best $10$ solutions from each side for each run of each variant, perform a round-robin tournament, and compute the ELO score (Elo,, 1978) (Fig. 5).
@@
The Coverage measures the volume of novelty discovered by GAME with the VEM. One limitation is that this behavior is dependent on two solutions. Another limitation is that visual behavior may not capture the full range of novelty. We use another measure similar to PATA-EC proposed in Enhanced-POET. The idea is that a BT is different from another if, in a tournament, the ranking it generates in the opposition is different. Using the intergenerational-tournament, we compute the ranking of all the BTs, and count the number of new rankings per generation compared to the previous generations (Fig. 3.e). The idea is that an open-ended system will continue to create new challenges that require new solutions.
@@ Results
Figure 5: Round robin tournament ELO score between the $10$ best solutions of each replication. GAME-SO is significantly better than all variants.
@@
As expected, GAME-SO and Diversity-only lead to the largest increases in BT size across generations. Diversity-only is the only variant showing a constant increase in size, suggesting that fitness maximization indirectly leads to pruning.
@@
GAME-SO also demonstrates the largest increase in behavioral complexity, suggesting that minimizing BT size prunes necessary stepping stones that lead to more complex behaviors. Diversity-only doesn’t show an increase in behavioral complexity similar to its BT size growth, showing that increased size doesn’t necessarily lead to greater behavioral complexity.
@@
Quality-only and GAME-MO (no VEM) exhibit significantly less diversity than other variants, which show similar diversity levels, with Diversity-only being slightly superior. Additionally, using the VEM results in less variance in coverage between replications, further validating its use as a behavior space.
Quality and Visual Diversity (Fig. 3.d)
Quality-only and GAME-MO (no VEM) show significantly worse QD-Scores (due to their poorer coverage). Diversity-only shows only a slightly worse QD-Score, suggesting that diversity alone is already a powerful optimizer, though the best QD-Score is achieved when also considering quality.
@@
Comparing the best solutions’ quality, we observe that GAME-SO produces significantly better BTs. This suggests that increased behavioral complexity leads to significant performance improvements and that directly attempting to reduce solution size with a secondary objective prevents the discovery of the highest-performing solutions. GAME-MO (no VEM) also produces high-performing solutions. One explanation is that the handcrafted behavior space correlates more with the fitness, thus focusing the search toward better-performing solutions.
@@
The GAME-SO and GAME-MO show similar open-endedness, which slightly decreases with each iteration, finding $80$% new tasks in the 2nd and 3rd generations and declining to $60$% by the final generations. GAME-MO (no bootstrap) consistently generates $90$% novel BTs, suggesting this variant is the only one showing signs of true open-endedness throughout all $20$ generations. The other variants demonstrate less open-endedness, with Quality-Only converging to 00% novelty after eight generations in one replication. This further demonstrates that diversity prevents getting trapped in local minima.
@@
Figure 6: Global trends: Percentage of elites using (a) the Attack atomic with different targets and (b) the Go_to atomic for different thresholds through the generations. The solid line represents the median, and the shaded area between the min and max of GAME-MO’s three replications.
@@
Examining the Go_to atomics, all replications favor using a threshold distance equal to $50$% of the sight range while avoiding exact positioning, i.e., 00% (Fig. 6.b). This likely represents a balance between spending time moving precisely to the target position (i.e., 00%), which prevents units from performing other actions, and merely moving within sight of the target (i.e., $100$%), which limits visibility of what can be observed from the target position itself.
@@
We found an “arms race” in one of GAME-SO’s replications (Fig 7). The Red elites increased their focus on the archers, leading to the Blue elites’ counter-strategy to focus their healing on their archers.
Figure 7: Arms race example: Percentage of Red elites using the Attack atomic on weakest enemy archer and Blue elites using the Heal atomic to random ally archers through the generations for GAME-SO’s three replications. For replication $2$, as more Red elites targeted the Blue archers, Blue elites evolved to heal their archers.
Discussion
@@
We hypothesized that minimizing the BT size as a secondary objective wouldn’t impact quality and lead to smaller, more interpretable BTs. However, the results indicate it prunes the neutral mutations that seem essential stepping stones for high-performing solutions. This supports the neutral theory of molecular evolution (Kimura,, 1979), which states that most variation at the molecular level is neutral yet leads to the evolution of complex organisms in nature.
@@
We also hypothesized that bootstrapping each generation using solutions from previous ones would accelerate the search compared to starting from scratch. The results confirm this hypothesis. Nonetheless, GAME (no bootstrap) is the only variant demonstrating constant generation of new solutions throughout all $20$ generations. This phenomenon could be related to extinction events (Lehman and Miikkulainen,, 2015) and warrants further investigation, as it currently doesn’t yield the best diversity or quality.

### v3

@@ Introduction
  • GAME-MO: full variant with a multi-objective fitness that first maximizes the opposing side’s depleted health and secondarily minimizes the size of the BT;
  • GAME-SO (no bootstrap): GAME-SO with an ablation of the bootstrapping between generations;
  • GAME-SO (no VEM): GAME-SO with a 2D handcrafted behavior space equal to the evaluated side’s average remaining health and the time before completion (i.e., one side wins or $100$ steps timeout) instead of a VEM;
  • Diversity-only: do not use fitness, only rule (e.1);
  • Quality-only: MT-ME instead of MTMB-ME, i.e., each task has only one solution.
We run three replications of $20$ generations for each variant with 100 000100\,000 evaluations of $100$ steps per generation, $100$ tasks per generation, and $25$ cells per task’s archive.
@@
After each run, we conduct a tournament between each generation’s tasks, $10$ Blue generations against $10$ Red generations (i.e., 1 0001\,000 Red elites against 1 0001\,000 Blue elites), resulting in 1 000 0001\,000\,000 evaluations. This allows us to study whether there are intergenerational improvements.
@@
To measure variations in solution size, we compute the number of nodes of the BT elites (Fig. 3.a), and for behavioral complexity, we measure the entropy of the distribution of actions performed over time and units of the same side (Fig. 3.b).
Visual Diversity and Quality-Weighted Visual Diversity
Coverage and QD-Score are classic measures of QD algorithms but are not easily computed using the archive in a large behavior space. As a proxy, we compute a single 2D projection using PCA (the two components capturing 2222% and 1212% of the explained variance) for all the behaviors of the intergenerational tournaments for all replications (Fig. 4), discretize this space with a $100\times 100$ grid, and compute the corresponding elites. Coverage (Fig. 3.c) corresponds to the proportion of non-empty cells, and QD-Score (Fig. 3.d) to the average fitness of the elites.
@@
One limitation of computing the quality directly from the fitness in the QD-Score is that the fitness can be high because the opposite side is weak, not because the winning side is strong. To measure a less ambiguous quality of the elites found by each method, we follow Dharna et al., (2024) and select the best $10$ solutions from each side for each run of each variant, perform a round-robin tournament, and compute the ELO score (Elo,, 1978) (Fig. 5).
@@
Coverage measures the volume of novelty discovered by GAME with the VEM. One limitation is that this behavior is dependent on two solutions. Another limitation is that visual behavior may not capture the full range of novelty. We use another measure similar to PATA-EC proposed in Enhanced-POET. The idea is that a BT is different from another if, in a tournament, the ranking it generates in the opposition is different. Using the intergenerational tournament, we compute the ranking of all the BTs and count the number of new rankings per generation compared to the previous generations (Fig. 3.e). The idea is that an open-ended system will continue to create new challenges that require new solutions.
@@ Results
Figure 5: Round-robin tournament ELO score between the $10$ best solutions of each replication. GAME-SO is significantly better than all variants.
@@
Quality-only has a significantly higher increase in BT size. This can be explained by the lack of diversity, which focuses the optimization on a smaller set of solutions (one per task instead of $N_{cells}=25$), thus allowing more mutations to accumulate. Diversity-only has the second-largest increase in BT size, which can be explained by the absence of fitness regulation, leading to preserving “bad” nodes that would have been removed otherwise. Among all the GAME variants, GAME-SO shows a slightly higher increase in BT size.
@@
Quality-only also has the highest increase in behavioral complexity, with the same explanation as before. What is more interesting is the presence of GAME-SO as a close second. Compared to GAME-MO, this suggests that minimizing BT size prunes necessary stepping stones that lead to more complex behaviors. Diversity-only is ranked fourth, showing that increased BT size does not necessarily lead to greater behavioral complexity.
@@
Quality-only and GAME-SO (no VEM) exhibit significantly less diversity than other variants, which show similar diversity levels, with Diversity-only being slightly superior, validating the use of a VEM as a behavior space.
Quality-Weighted Visual Diversity (Fig. 3.d)
Quality-only and GAME-SO (no VEM) show significantly worse QD-Scores (due to their poorer coverage). Diversity-only shows only a slightly worse QD-Score, suggesting that diversity alone is already a powerful optimizer, though the best QD-Score is achieved when also considering quality.
@@
Comparing the best solutions’ quality, we observe that GAME-SO and some instances of Quality-only produce significantly better BTs. This suggests that increased behavioral complexity leads to significant performance improvements and that directly attempting to reduce solution size with a secondary objective may prevent the discovery of the highest-performing solutions. Quality-only has the largest variance between replications; this result suggests that diversity, quality, the VEM, and the bootstrap phase are all necessary to discover the best BTs.
@@
GAME-SO and GAME-MO show similar open-endedness, which slightly decreases with each iteration, finding 8080% new tasks in the 2nd and 3rd generations and declining to 6060% by the final generations. GAME-SO (no bootstrap) consistently generates 9090% novel BTs, suggesting this variant is the only one showing signs of true open-endedness throughout all $20$ generations, but also that bootstrapping may be detrimental to open-endedness. The other variants demonstrate less open-endedness, with Quality-only getting close to 0% novelty. This further demonstrates that diversity prevents getting trapped in local minima.
@@
Figure 6: Global trends: Percentage of elites using (a) the Attack atomic with different targets and (b) the Go_to atomic for different thresholds through the generations. The solid line represents the median, and the shaded area shows the min and max of GAME-MO’s three replications.
Figure 7: Arms race example: Percentage of Red elites using the Attack atomic on the weakest enemy archer and Blue elites using the Heal atomic on random ally archers through the generations for GAME-SO’s three replications. For replication 22, as more Red elites targeted the Blue archers, Blue elites evolved to heal their archers.
@@
Examining the Go_to atomics, all replications favor using a threshold distance equal to 5050% of the sight range while avoiding exact positioning, i.e., 0% (Fig. 6.b). This likely represents a balance between spending time moving precisely to the target position (i.e., 0%), which prevents units from performing other actions, and merely moving within sight of the target (i.e., 100100%), which limits visibility of what can be observed from the target position itself.
@@
We found an arms race in one of GAME-SO’s replications (Fig. 7). The Red elites increased their focus on the archers, leading to the Blue elites’ counter-strategy to focus their healing on their archers.
Case study: 2D Soft robots, Wrestling
Environment
Replicating the biological evolution feat of creating the diversity of living creatures’ morphology has been a long-standing research topic in the artificial life community (Langton,, 1997). One early work involves evolving articulated rigid-body creatures in an adversarial problem where the goal is for each creature to get as close as possible to a central cube (Sims,, 1994). To move closer to real creatures, some more recent works focus on soft robots with the problem of locomotion (Cheney et al.,, 2014; Kriegman et al.,, 2017; Cheney et al.,, 2018; Bhatia et al.,, 2021; Mertan and Cheney,, 2023). EvoGym (Bhatia et al.,, 2021) was proposed to facilitate the creation of environments for coevolving morphology and control of 2D soft robots.
While the coevolution of morphology and its control is an interesting and challenging topic, it is beyond the scope of this introductory GAME paper to explore the intersection of adversarial coevolution and body–brain coevolution. We instead focus on the adversarial coevolution of morphology in an adversarial environment, where the morphology passively provides actuation. We build a custom environment using EvoGym (Bhatia et al.,, 2021), which we refer to as Wrestling for 2D soft-robot simulation. We follow Cheney et al., (2014); Kriegman et al., (2017) by defining passive and active voxels that change surface following a sine wave pattern. We defined seven types of voxel: Empty (invisible); rigid (black); soft (gray); horizontal actuated in-phase (orange); vertical actuated in-phase (royalblue); horizontal actuated anti-phase (gold); vertical actuated anti-phase (skyblue).
We set the sine wave period to $12$ timesteps because it empirically results in most of the randomly sampled robots exhibiting non-stationary behaviors. The solution space is $5\times 5$ 2D soft robots, with the constraints that there is at least one actuated voxel and that all voxels are connected. See Fig. 8 for examples of different morphologies. The variation operator randomly chooses between adding, deleting, or mutating a voxel ($\frac{1}{3}$ for each) while respecting the constraints and does this $k$ times ($k=3$ in this paper).
In Wrestling, two robots start symmetrically on the edge of an arena of width 30 voxels. The fitness function is the percentage of timesteps the robot was closest to the center after $200$ timesteps. This leads each robot to reach the center as quickly as possible, push the other, and resist the other’s pushes. One specificity of this fitness function compared to the one used in Parabellum is that it is a relative measure of quality, i.e., given two competing blue and red solutions, $s_{blue}$ and $s_{red}$, $f(s_{blue})=1-f(s_{red})$. This means that even trivially bad solutions (e.g., standing morphologies) would lead to at least one of the two opposing fitnesses being positive. This prevents a direct evaluation of quality from the QD-Score and instead requires the use of a tournament or another measure of quality (e.g., the velocity of the solutions). This is interesting because a one-sided illumination would require already knowing a high-quality solution to avoid trivially overfitting a bad solution.
The behavior space is, similarly to Parabellum, the concatenation of the visual embedding of CLIP Radford et al., (2021) for five frames of the video. For speed optimization, we do not use EvoGym’s visualization but a reimplementation in JAX that directly creates RGB arrays for CLIP.
Results
Figure 8: Wrestling illumination with sampled snapshots. The central point cloud is the 2D PCA of the visual embedding of one illumination from GAME of the Wrestling adversarial problem (capturing 13.4% and 8.4% of the variance). The outer edge snapshots show examples of frames captured at the end of a diverse set of evaluations. The PCA captures slow-moving robots on the right and faster-moving robots on the left, each with a large variety of morphologies. See Figure_8.mp4 in the repository for a video of this figure.
Figure 9: Wrestling 2D PCA of the visual embedding with fitness. We concatenated all visual embedding behaviors and performed one PCA capturing 13.4% and 8.4% of the variance, and show the corresponding projection for GAME and Random (three replications). GAME leads to wider coverage than using the EvoGym random sampler. Wrestling’s fitness function measures the relative quality of solutions, so examining it does not directly allow comparison of the quality of the discovered solutions.
Figure 10: Wrestling visual diversity through generations. GAME leads to higher coverage of the visual behavior space. The solid line represents the median, and the shaded area shows the min and max of GAME and Random (three replications).
Figure 11: Wrestling quality of the best morphologies. ELO score computed with a tournament between the $10$ best morphologies of each side of each variant’s replications. GAME finds better morphologies than Random.
Comparison with a Random ablation.
To demonstrate the generality of GAME, we apply the same implementation to Wrestling as the one used on Parabellum with just a smaller budget of evaluations ($N_{budget}=$20\,000$\$, $N_{gen}=10$, and $N_{cells}=20$). We compare GAME against an ablation, Random, that uses EvoGym’s robot sampling function instead of applying the variation operator to an elite of the current generation. We performed $3$ replications for each and the same number of overall evaluations.
Fig. 9 shows the 2D PCA (capturing 13.4% and 8.4% of the variance) of the visual embedding. GAME leads to a larger coverage with a median of 44.2% [44.6, 39.7] against 28.9% [29.9, 27.3] for Random (Fig. 10) and higher quality when comparing the top-10 elites of each replication in a tournament; the lowest GAME ranking is higher than the highest Random ranking (Fig. 11). Fig. 9 does not allow us to compare the quality of the solutions due to Wrestling’s relative fitness function. To investigate the reason for the higher quality in the tournament, Fig. 12 shows the same PCA projection with the average velocity of the two soft robots. We can see that GAME leads to finding morphologies with higher velocities than the Random ablation. The main component of the PCA of the visual embedding captures slower robots on the right side, which can be visualized in Fig. 8 (and the supplementary video), with faster robots on the left side and slower-moving robots on the right side that do not reach the center of the arena.
Figure 12: Wrestling 2D PCA of the visual embedding with velocity. Same PCA projection but showing the average velocity of the two soft robots. GAME leads to discovering robot morphologies with higher velocities.
Lineage visualization.
GAME keeps track of each ancestor of a solution. Fig. 13 shows the full lineage of the best Red and Blue solutions of the inter-variant tournament. Because we mutate at most $k=3$, the solutions change smoothly. In comparison, using a higher $k$ or a crossover could lead to abrupt changes. Still, the initial and final morphologies have very few voxels in common. Future work could focus on using an indirect encoding (Cheney et al.,, 2014) to allow a crossover and boost the variation operator’s ability to explore the search space.
Figure 13: Ancestry of the best morphologies in the Wrestling tournament. GAME keeps track of each solution’s genealogy through the generations and evaluations. (a) Red. (b) Blue. The variation operator mutates at most $k=3$ voxels, leading to smooth transitions from one morphology to another.
Morphological speciation.
Figure 14: Morphological species through generations. (a–b) Clustering of the different morphologies into species from all generations using k-modes and UMAP for visualization. The numbers point to the selected elite of each cluster for each generation. (c–d) Average ELO score of each species through generations. GAME discovers all species in the first generation and then improves their average fitness.
To study the evolution of different morphologies through generations, we perform a posteriori morphological speciation for one of GAME’s replications. We first aggregate all the elites’ morphologies from each generation and perform clustering using k-modes ($k=5$) to identify different morphological species. Fig. 14.a–b show the UMAP projection of the clustering. Then, for each cluster at each generation, we compute the average ELO score in an intergenerational tournament. Fig. 14.c–d show the variation of performance through the generations for both sides, with a snapshot of the best morphology for each cluster and generation.
First, all the species are discovered in the first generation, even though the clustering is performed a posteriori on all generations. This means that GAME does not discover a significant new morphology after the first generation. This can be explained by two factors: (1) GAME is a QD algorithm, so it searches for diversity from the start, and (2) this specific problem is not open-ended.
Second, for most morphological species, the average ELO score improves over generations, demonstrating GAME’s ability to pressure for improved solution quality. We can notice that one Blue species disappears in the last generation. This discrepancy can be attributed to the difference between morphological diversity and the visual embedding diversity optimized by GAME, which captures other elements, such as velocity (due to the use of five frames), and the interaction between the two soft robots.
Case study: Deck building, Hearthbreaker
Figure 15: Hearthbreaker intergenerational tournament. ELO score of the elites from GAME execution on the five pairs of classes in Hearthbreaker in an intergenerational tournament for each pair. GAME leads to an overall improvement of the elites’ quality at each generation for both sides and all five pairs of classes.
Figure 16: Hearthbreaker archive comparison between GAME and ME. GAME leads to better top-50 elites than ME for 7 out of 10 comparisons and is not significantly better for the other 3. (Significance levels of the Mann–Whitney U test: $^{\ast}=p<0.05$, $^{\ast\ast}=p<0.01$, $^{\ast\ast\ast}=p<0.001$, and ns = not significant.)
Figure 17: Hearthbreaker measures comparison between GAME and ME. ME leads to denser and larger coverage due to focusing on only one task, while GAME illuminates 50 tasks simultaneously. Still, GAME finds solutions with similar quality to ME, even though it was not explicitly searching for solutions against the starting decks.
Warrior VS Warlock Rogue VS Paladin Hunter VS Druid Shaman VS Mage Mage VS Priest
(a) (b) (c) (d) (e) (f) (g) (h) (i) (j)
Coverage GAME 33.8% 34.3% 32.3% 33.5% 35.5% 30.0% 36.3% 27.3% 38.3% 33.0%
ME 66.0% 70.8% 65.3% 61.5% 66.3% 69.3% 64.0% 63.8% 66.8% 71.5%
QD-Score GAME 6.3 4.7 7.1 7.1 5.6 3.4 8.2 5.8 5.6 7.5
ME 12.5 9.1 13.3 13.1 10.8 6.2 14.9 12.1 9.3 16.4
Best Elite GAME 24.6 23.7 27.6 27.3 27.1 21.8 26.6 27.5 24.1 27.2
ME 26.1 23.4 27.9 27.1 26.7 22.6 28 28.0 25.3 28.2
Table 1: Illumination comparison between GAME and ME in Hearthbreaker against the starting deck. Coverage, QD-Score, and best-elite fitness of each execution of GAME and ME against the corresponding starting deck (Fig. 17). ME leads to higher coverage and QD-Score but similar highest fitness.
Environment
As a third adversarial problem, we aimed to compare GAME with a classic MAP-Elites approach that fixes one side of the adversarial problem, as in Fontaine et al., (2019). We use a Python simulator of the Hearthstone digital collectible card game (Blizzard Entertainment,, 2014), Hearthbreaker Daniel et al., (2014), for easy interfacing with our implementation of GAME.
Hearthstone is a two-phase card game. In the first phase, deck building, the player chooses a class (among the $9$ available) and creates a deck of $30$ cards. Then, in the second phase, deck playing, the player picks their deck of cards and plays a game against another player with their own class and deck of cards. The goal is to set the opponent’s life from $30$ to 0. In this case study, similarly to Fontaine et al., (2019), we only consider the deck building phase and use the Hearthbreaker Trade Agent heuristic to play the deck. This heuristic is a greedy, handcrafted policy that maximizes the loss of minions for the opponent and subsequently maximizes the damage to the opponent’s health.
We attempted to follow the method described in the original paper as closely as possible and did not use the VEM contribution of this paper as the behavior space. The solution space consists of a set of $30$ cards, comprising both class-specific cards and common cards (with between 250 and 300 available cards for each class in Hearthbreaker), with no more than two copies of the same card. Available cards may differ between the different versions of Hearthstone. We used the same variation operator, i.e., we pick one elite at random, pick $k$ from a geometric distribution with $P(k=1)=0.5$, and then randomly replace $k$ cards from the elite deck using a uniform distribution over the valid cards. The behavior space is two-dimensional: the deck’s mana cost mean and standard deviation. We use the same unstructured archive, as it can handle both small and large spaces without the same value of hyperparameter, i.e., the number of cells in the archive. The fitness function is the difference in life between the player and the opponent. As the game ends when one player reaches 0, the fitness function is positive when the player wins (and is bounded by the starting health of $30$) and negative when the player loses (and is bounded by 30-30). As the game is stochastic by nature due to the shuffling of the cards, the fitness is averaged over $n=50$ duels with fixed seeds, with half of them starting with the player and the other half with the opponent.
Results
We compare GAME against an ablation, ME, corresponding to MAP-Elites, i.e., only one generation and one task, with the same total budget of 40 00040\,000 evaluations per side and number of cells in the archive. GAME is set with $N_{gen}=8$, $N_{budget}=$10\,000$\$, $N_{tasks}=50$, and $N_{cells}=$20$\$ and ME with $N_{gen}=1$, $N_{budget}=$40\,000$\$, $N_{tasks}=1$, and $N_{cells}=$1\,000$\$. We chose five pairs of classes that are empirically known to be balanced: Warrior and Warlock, Rogue and Paladin, Hunter and Druid, Shaman and Mage, and Mage and Priest. ME independently evolves the deck of one side against the fixed starting deck of the other side (similarly to Fontaine et al., (2019)) and is thus executed once for each side. In contrast, GAME coevolves both decks in one run. Fig. 15 shows the intergenerational ELO score of the five GAME executions on each pair of classes. We can see that for all instances, the elite’s average quality improves at each generation.
As GAME and ME do not use the same resolution for the archive, for a fairer comparison, we recomputed both variants’ archives in a 2D grid with a bin size of $0.1$, and performed a tournament between the new archives’ elites for each pair of classes. Fig. 16 shows the resulting ELO scores. We can see that GAME finds significantly better solutions for 7 out of 10 comparisons and non-significantly better solutions for the remaining three. A simple reason is that ME decks are only evolved to compete against the starting deck, whereas GAME coevolves a diversity of decks that prevent overfitting to a single one.
To compare the illumination, we evaluated each elite from GAME against the corresponding starting deck (the same used as a fixed opponent for ME) and compared the corresponding diversity and quality. Fig. 17 shows each archive, and Tab. 1 shows the corresponding Coverage, QD-Score, and best-elite fitness. GAME’s coverage is significantly worse than ME’s. The reason is that it splits the 1 0001\,000 cells into $50$ tasks, thus leaving only $20$ cells for diversity, while ME can allocate the 1 0001\,000 cells for diversity. This becomes apparent in this case study because we use a much smaller 2D behavior space, rather than a visual embedding. It is thus much easier to cover the entire reachable behavior space. One limitation of GAME is the independence of diversity in each task, which leads to storing elites with similar behaviors across multiple tasks, thereby reducing the overall diversity of the archive. Still, even though GAME’s elites are not evolved against the starting decks like those from ME, they show similar performance against them.
Discussion and Future Work
Multi-agent battle game: Parabellum
@@
We hypothesized that minimizing the BT size as a secondary objective would not impact quality and would lead to smaller, more interpretable BTs. However, the results indicate it prunes the neutral mutations that seem to be essential stepping stones for high-performing solutions. This supports the neutral theory of molecular evolution (Kimura,, 1979), which states that most variation at the molecular level is neutral yet leads to the evolution of complex organisms in nature.
@@
We also hypothesized that bootstrapping each generation using solutions from previous ones would accelerate the search compared to starting from scratch. The results confirm this hypothesis. Nonetheless, GAME-SO (no bootstrap) is the only variant demonstrating constant generation of new solutions throughout all $20$ generations. This phenomenon could be related to extinction events (Lehman and Miikkulainen,, 2015) and warrants further investigation, as it currently does not yield the best diversity or quality.
@@
One limitation of this case study is the use of BTs as controllers. They are inherently interpretable but require handcrafted atomics that limit the possible behaviors. An interesting future direction is to utilize end-to-end neural networks that map observations to actions, potentially employing neuroevolution Stanley and Miikkulainen, (2002), to move toward a more open-ended search space.
2D Soft robots: Wrestling
The main limitation of this case study is the use of passive actuation to control the robots, which, in particular, prevents the robots from exhibiting reactive behavior based on sensor inputs, a key component of intelligent behaviors (Brooks,, 1991; Pfeifer and Bongard,, 2006). Future work should focus on incorporating body–brain coevolution with GAME to move toward a truly open-ended search space for artificial creatures.
Deck building: Hearthbreaker
There are two limitations of the current approach. First, the behavior space may be too simple and is independent of the duels (as it is purely computed from the deck and not influenced by how the deck is played), which limits the possible diversity of decks. Second, we use a greedy heuristic that ignores the combo aspect of the game and focuses only on building decks with many minions (all decks evolved with GAME have at least 86% (26 out of 30) minion cards, with a median of 97% (29 out of 30). In contrast, minions correspond to 65% of the available cards, which would correspond to 20 out of 30 cards in the deck. Future work should focus on using a more advanced agent, such as MCTS (Santos et al.,, 2017), to explore the game’s combo aspect and thus significantly increase the open-endedness of this adversarial problem.